1. Calculate what mass of water can be changed to steam if 10.6kJ of heat energy is supplied to the water at 1000C. (Specific latent heat of vaporization of water = 2.26x106J/kg)




  1. Ammonia is vaporised in order to freeze an ice rink.

a)  Find out how much heat it would take to vaporise 1 g of ammonia.

    (Specific latent heat of vaporisation of ammonia = 1.34 x 10J/kg      Specific latent heat of fusion of ice = 3.34 x 10J/kg).






1
Expert's answer
2021-06-07T19:16:11-0400

1)

"Q=Lm,\implies m=\frac QL,"

"m=\frac{10.6\cdot 10^3}{2.26\cdot 10^6}=4.7~g,"

2)

"Q=Lm,"

"Q=1.34\cdot 10^6\cdot 1\cdot 10^{-3}=1.34~kJ"